# Sharp Analysis for KL-Regularized Contextual Bandits and RLHF

### v1

@@ Abstract
Reverse-Kullback-Leibler (KL) regularization has emerged to be a predominant technique used to enhance policy optimization in reinforcement learning (RL) and reinforcement learning from human feedback (RLHF), which forces the learned policy to stay close to a reference policy. While the effectiveness and necessity of KL-regularization have been empirically demonstrated in various practical scenarios, current theoretical analysis of KL-regularized RLHF still obtains the same $\mathcal{O}(1/\epsilon^{2})$ sample complexity as problems without KL-regularization. To understand the fundamental distinction between policy learning objectives with KL-regularization and ones without KL-regularization, we are the first to theoretically demonstrate the power of KL-regularization by providing a sharp analysis for KL-regularized contextual bandits and RLHF, revealing an $\mathcal{O}(1/\epsilon)$ sample complexity when $\epsilon$ is sufficiently small.
We further explore the role of data coverage in contextual bandits and RLHF. While the coverage assumption is commonly employed in offline RLHF to link the samples from the reference policy to the optimal policy, often at the cost of a multiplicative dependence on the coverage coefficient, its impact on the sample complexity of online RLHF remains unclear. Previous theoretical analyses of online RLHF typically require explicit exploration and additional structural assumptions on the reward function class. In contrast, we show that with sufficient coverage from the reference policy, a simple two-stage mixed sampling strategy can achieve a sample complexity with only an additive dependence on the coverage coefficient. Our results provide a comprehensive understanding of the roles of KL-regularization and data coverage in RLHF, shedding light on the design of more efficient RLHF algorithms.
@@ 1 Introduction
Recently, Reinforcement Learning from Human Feedback (RLHF) has emerged as a central tool for aligning large language models (LLMs) and diffusion models with human values and preferences (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022; Rafailov et al., 2024), exhibiting impressive capabilities in applications, such as Chatgpt (Achiam et al., 2023), Claude (Anthropic, 2023), Gemini (Team et al., 2023), and LLaMA-3 (Meta, 2024).
@@
RLHF methods treat the language model as a policy that takes a prompt $x$ and produces a response $a$ conditioned on $x$, and they optimize the policy by aligning it with human feedback. There are mainly two kinds of feedback: absolute rating and preference comparison. For absolute rating, the collection typically involves human annotators to provide rating scores like 1 to 5 (Wang et al., 2024a, b) for responses or hard 0-1 scores for math reasoning tasks since the reasoning tasks often have gold standard answers (Cobbe et al., 2021; Hendrycks et al., 2021; Xiong et al., 2024b).
On the other hand, preference comparison is frequently applied in chat tasks when making comparisons is much easier for human labeler (Achiam et al., 2023). Although preference feedback is believed to be more intuitive for human users and easier to collect, it also poses more challenges for the RLHF algorithms to effectively leverage the feedback signals since the reward signals are not directly observed. In practice, the learning process typically involves (a) constructing a reward model based on the maximum likelihood estimation (MLE) of Bradley-Terry (BT) (Bradley and Terry, 1952a) model from the preference feedback; (b) applying RL algorithms like PPO (Schulman et al., 2017b) to train the language model so that it maximizes the reward signals with KL regularization (Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023). There is also a series of work on direct preference learning algorithms, including Slic (Zhao et al., 2023b), DPO (Rafailov et al., 2024), IPO (Azar et al., 2024), which directly optimize the preference signals to train the language model without constructing the reward model. However, despite their efficiency, these types of algorithms often suffer from overfitting, and DPO suffers from a drop of chosen probability (Song et al., 2024; Liu et al., 2024; Amini et al., 2024; Huang et al., 2024).
@@
Since human value and preference are so complicated that they are unlikely to be encompassed by the considered preference model classes, the learned model is easy to be hacked and become biased. Practically, the policy may generate disproportionate bold words or emoji to please the learned reward (Zhang et al., 2024). Hence, the KL-regularization between the learned policy and a reference policy (the pre-trained model after supervised fine-tuning) plays a fundamental role in RLHF to avoid overfitting. There is a line of RLHF work that realizes the significance of KL-regularization and regards the problem as a reverse-KL regularized contextual bandit (Ziegler et al., 2019; Wu et al., 2021; Ouyang et al., 2022; Rafailov et al., 2024; Xiong et al., 2024a; Ye et al., 2024b). However, they adopt the techniques from bandit framework and neglect the characteristic of reverse-KL-regularization, thus obtaining almost the same sample complexity with problems without KL-regularization. Therefore, the question of whether there exists a fundamental distinction between policy learning objectives with and without KL-regularization is still largely under-explored.
@@
Compared to the offline RLHF algorithms (Rafailov et al., 2024; Azar et al., 2024; Chen et al., 2024) that can only use planning to approximate the solution to the relative entropy minimization problem (Ziebart et al., 2008; Song et al., 2024), online RLHF has been demonstrated to outperform offline methods empirically and theoretically (Bai et al., 2023; Meta, 2024; Xiong et al., 2024a; Tajwar et al., 2024; Song et al., 2024; Wu et al., 2024b), because it has further interactions with human or the preference oracle. Most standard theoretical online RL techniques apply optimism to balance exploration and exploitation (Abbasi-Yadkori et al., 2011; Wang et al., 2020; Jin et al., 2021). However, it is inefficient to implement exploration for practical RLHF algorithms. Meanwhile, an emerging line of offline RLHF literature highlights the coverage of the reference policy $\pi_{0}$. The coverage of $\pi_{0}$ refers to the ability of the model to generate diverse responses for a wide range of prompts. A model with good coverage can generalize well to unseen contexts and actions, which is essential for the learned reward function to also generalize well. In practice, this is evidenced by the fact that the simple best-of-$n$ sampling based on $\pi_{0}$ is competitive with the well-tuned PPO algorithm for general open-ended conversation tasks (Dong et al., 2023), and the fact that the $\pi_{0}$ can solve a majority of the math problems with multiple responses (Shao et al., 2024; Nakano et al., 2021). However, the theoretical understanding of the role of coverage in online RLHF is still largely understudied. Thus, it is natural to ask if explicit exploration is necessary for online RLHF with good coverage of $\pi_{0}$ and how the coverage of $\pi_{0}$ affects the sample complexity of online RLHF.
In this paper, we answer the above questions by
  • providing a novel fine-grained analysis for KL-regularized contextual bandits and RLHF, which adapts to the optimization landscape of the reverse-KL regularization and reveals a sharper sample complexity than the existing results,
  • proposing an efficient 2-stage mixed sampling strategy for online RLHF with good coverage of $\pi_{0}$, which achieves sample complexity with only an additive dependence on the coverage coefficient.
@@ 1.1 Our Contributions
In this work, we make a first attempt to illustrate the statistical benefits of KL-regularization for policy optimization in contextual bandits and reinforcement learning (RL) from preference feedback. Our main contributions are summarized as follows:
  • In Section 3, we formulate RLHF with absolute-rating feedback as a contextual bandit problem with KL-regularization. First, we provide a lower bound for the KL-regularized contextual bandit problem, which indicates that the sample complexity of the problem is $\Omega(\eta\log N_{\mathcal{R}}(\epsilon)/\epsilon)$ when $\epsilon$ is sufficiently small, where $N_{\mathcal{R}}(\epsilon)$ is the covering number of the reward function class and $\eta$ is the KL-regularization coefficient.
  • Then we showcase a novel analysis to upper bound the suboptimality gap of the KL-regularized objective in contextual bandits, and propose a simple two-stage mixed sampling strategy for online RLHF which achieves a sample complexity of $\mathcal{O}(\max(\eta^{2}D^{2},\eta/\epsilon)\log N_{\mathcal{R}}(\epsilon/\delta))$ when the reward scale is a constant, where $D$ is the coverage coefficient of the reference policy $\pi_{0}$ and $\delta$ is the confidence parameter. To the best of our knowledge, this is the first work to provide a sharp sample complexity for KL-regularized contextual bandits.
  • In Section 4, we extend our analysis to reinforcement learning from preference feedback. We rigorously demonstrate that KL-regularization is essential for more efficient policy learning in RLHF with preference data. We further propose a two-stage mixed sampling strategy for online preference learning setting with good coverage of $\pi_{0}$, which achieves a sample complexity of $\mathcal{O}(\max(\eta^{2}D^{2},\eta/\epsilon)\log N_{\mathcal{R}}(\epsilon/\delta))$ when the reward scale is a constant.
@@ 1.2 Previous Understanding of KL Regularization in RL
While we mainly focus on the theoretical understanding of KL-regularization in RLHF, it is also worth mentioning that our analysis for KL-regularized contextual bandits also contributes to the theoretical understanding of the impact of KL-regularization in RL since contextual bandits can be viewed as a simplified version of Markov decision processes (MDPs).
In RL, KL-regularization has been widely used to stabilize the learning process and prevent the policy from deviating too far from the reference policy. Here, we provide a brief overview of the existing understanding of KL-regularization in decision-making problems. From the perspective of policy optimization, KL-regularization captures entropy regularization as a special case  [^1] We can regard the entropy regularization as a special case of KL-regularization by setting the reference policy as the uniform distribution., which is also an extensively used technique in RL literature (Sutton, 2018; Szepesvári, 2022). There is a large body of literature that has explored the benefits of entropy regularization or KL-regularization in RL (Schulman et al., 2015; Fox et al., 2016; Schulman et al., 2017a; Haarnoja et al., 2017, 2018; Ahmed et al., 2019). Most related to our work, Ahmed et al. (2019) provided a comprehensive understanding of the role of entropy regularization in RL, showing that entropy regularization can improve the training efficiency and stability of the policy optimization process by changing the optimization landscape through experiments on continuous control tasks (Brockman, 2016).
Theoretically, Neu et al. (2017) provided a unified view of entropy regularization as approximate variants of Mirror Descent or Dual Averaging, and left the statistical justification for using entropy regularization in RL as an open question. Geist et al. (2019) provided a framework for analyzing the error propagation in regularized MDPs, which also focused on the proof of the convergence for the policy optimization methods with regularization and lacked a sharp sample complexity analysis.
1.3 Other Related Work
Analyses for Policy Optimization with Regularization
While it is previously unknown whether regularization can improve the sample complexity of policy optimization without additional assumptions, there are some works that provided a sharp convergence rate in the presence of regularization (Mei et al., 2020; Shani et al., 2020; Agarwal et al., 2020, 2021). However, these works either assumed the access of exact or unbiased policy gradient or required uniform value function approximation error, which are not the standard case in sample-based RL setting.
@@ 1.3 Other Related Work
There are mainly three types of RLHF algorithms: offline, online and hyrbid. The most well-known offline algorithms are Slic (Zhao et al., 2023b), Direct Preference Optimization (DPO) (Rafailov et al., 2024), Identity-PO (IPO) (Azar et al., 2024) and (SPIN) (Chen et al., 2024). They aim to approximate the closed-form solution of the optimization problem on a fixed offline dataset. For the online algorithms, the most representative one is Proximal Policy Optimization (PPO) (Schulman et al., 2017b). PPO has been used in the Chat-GPT (OpenAI, 2023), Gemini (Team et al., 2023), and Claude (Bai et al., 2022). However, the deep RL method PPO is known to be sample inefficient and unstable, making its success hard to reproduce for the open-source community. In response to this, there have been many efforts to propose alternative algorithms to the PPO algorithm. The Reward ranked fine-tuning (RAFT) (also known as rejection sampling finetuning) (Dong et al., 2023; Touvron et al., 2023; Gulcehre et al., 2023; Gui et al., 2024) is a stable framework requiring minimal hyper-parameter tuning, which iteratively learns from the best-of-n policy (Nakano et al., 2021). This framework proves to be particularly effective in the reasoning task such as (Gou et al., 2024; Tong et al., 2024). However, the RAFT-like algorithms only use the positive signal by imitating the best-of-n sampling. To further improve the efficiency, there is an emerging body of literature that proposes online direct preference optimization by extending DPO or IPO to an online iterative framework (Xiong et al., 2024a; Guo et al., 2024; Wu et al., 2024b; Calandriello et al., 2024; Xiong et al., 2024b). Finally, for the third type, the common point of hybrid and online algorithms is that they both require further interaction with the preference oracle and on-policy data collection. The difference is that hybrid algorithms start with a pre-collected dataset (Xiong et al., 2024a; Song et al., 2024; Touvron et al., 2023), while the online algorithms learn from scratch.
@@
The theoretical study of RLHF can date back to the dueling bandits (Yue et al., 2012; Dudík et al., 2015; Saha, 2021; Di et al., 2024; Li et al., 2024; Verma et al., 2024; Wu et al., 2024a) and follow-up work on MDPs (Wang et al., 2023a; Zhu et al., 2023). However, these works deviate from the practice because they do not realize the significance of KL-regularization and still choose the greedy policy that simply maximizes the reward. After this line of work, Xiong et al. (2024a); Ye et al. (2024b); Song et al. (2024) highlight the KL-regularization theoretically and incorporate the KL term into the learning objective. However, they circumvent the special advantages of KL-regularization and still follow the techniques in bandit analysis, thus obtaining loose bounds. In our paper, we establish a new lower bound and a sharper upper bound for the KL-regularized framework, thus validating the empirical advantage of KL-regularization. There are also some works extending KL-regularized RLHF from bandit problems to the Markov decision process (MDP) problems (Zhong et al., 2024; Xiong et al., 2024b). We expect that our techniques can also be extended to the MDP setting, which we leave for future work.
2 Preliminaries
In this section, we formally state the problem settings of RL from human feedback (RLHF), where we consider two types of feedback: absolute rating and preference.
2.1 Contextual Bandits with KL Regularization
The first setting is the absolute-rating feedback, where we can query the ground-truth reward function to measure the quality of the responses by providing absolute reward value. For instance, in the NVIDIA Helpsteer project (Wang et al., 2023b, 2024c), human labelers are required to provide absolute score in five attributes: helpfulness, correctness, coherence, complexity, and verbosity. The dataset leads to many high-ranking open-source reward models, including the ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024a, b), URM-LLaMa-3.1-8B [^2] https://huggingface.co/LxzGordon/URM-LLaMa-3.1-8B, and Llama-3.1-Nemotron-70B-Reward [^3] https://huggingface.co/nvidia/Llama-3.1-Nemotron-70B-Reward. We also notice that recently this feedback framework is extended to other task such as video generation (He et al., 2024).
The absolute-rating feedback is directly modeled as reward functions (Wang et al., 2024a; Xiong et al., 2024b), and is regarded as contextual bandits with KL regularization. In the contextual bandit setting, at each round $t$, the agent observes a context $x_{t}\in\mathcal{X}$ generated from a distribution $d_{0}$ and chooses an action $a_{t}\in\mathcal{A}$. The agent receives a stochastic reward $r_{t}\in\mathbb{R}$ depending on the context $x_{t}$ and the action $a_{t}$. The goal is to maximize the expected cumulative reward over $T$ rounds.

### v2

@@ Abstract
Reverse-Kullback-Leibler (KL) regularization has emerged to be a predominant technique to enhance policy optimization in reinforcement learning (RL) and reinforcement learning from human feedback (RLHF), which forces the learned policy to stay close to a reference policy. While the effectiveness of KL-regularization has been empirically demonstrated in various practical scenarios, current theoretical analyses of KL-regularized RLHF still yield the same $\mathcal{O}(1/\epsilon^{2})$ sample complexity as ones without KL-regularization. To understand the fundamental distinction between objectives with KL-regularization and ones without KL-regularization, we are the first to theoretically demonstrate the power of KL-regularization by providing a sharp analysis for KL-regularized contextual bandits and RLHF, revealing an $\mathcal{O}(1/\epsilon)$ sample complexity when $\epsilon$ is sufficiently small. We also prove matching lower bounds for both settings. More specifically, we study how the coverage of the reference policy affects the sample complexity of KL-regularized online contextual bandits and RLHF. We show that with sufficient coverage from the reference policy, a simple two-stage mixed sampling algorithm can achieve an $\mathcal{O}(1/\epsilon)$ sample complexity with only an additive dependence on the coverage coefficient, thus proving the benefits of online data even without explicit exploration. Our results provide a comprehensive understanding of the roles of KL-regularization and data coverage in online decision making, shedding light on the design of more efficient algorithms.
@@ 1 Introduction
The KL-regularized contextual bandit problem (Langford and Zhang, 2007; Xiong et al., 2024a) has raised tremendous interest recently because of the significant development of the post-training stage in large language models (LLMs) and diffusion models from preference feedback (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022; Rafailov et al., 2024), which is called Reinforcement Learning from Human Feedback (RLHF). RLHF aims to optimize the policy by aligning it with human feedback, exhibiting impressive capabilities in applications such as Chatgpt (Achiam et al., 2023), Claude (Anthropic, 2023), Gemini (Team et al., 2023), and LLaMA-3 (Meta, 2024).
@@
In RLHF, we treat the language model as a policy that takes a prompt $x$ and produces a response $a$ conditioned on $x$, optimizing the policy by aligning it with human feedback. There are mainly two kinds of feedback: absolute rating and preference comparison. For absolute rating, the collection typically involves human annotators to provide rating scores like 1 to 5 (Wang et al., 2024a, b) for responses or hard 0-1 scores for math reasoning tasks since the reasoning tasks often have gold standard answers (Cobbe et al., 2021; Hendrycks et al., 2021; Xiong et al., 2024b). Although discrete feedback is believed to be more intuitive for human users and easier to collect, it also poses more challenges for the RLHF algorithms to effectively leverage the feedback signals since the reward signals are not directly observed. In practice, the learning process typically involves (a) constructing a reward model based on the maximum likelihood estimation (MLE) of Bradley-Terry (BT) (Bradley and Terry, 1952a) model from the preference feedback; (b) applying RL algorithms like PPO (Schulman et al., 2017b) to train the language model so that it maximizes the reward signals with KL regularization (Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023).
@@
Since the human feedback data only covers a tiny fraction of possible interactions, if we optimize the model purely for the reward without constraints, it might learn behaviors that work well for the training feedback but fail catastrophically on slightly different inputs. For example, the policy may generate disproportionate bold words or emoji to please the learned reward (Zhang et al., 2024). Hence, the KL-regularization between the learned policy and a reference policy (the pre-trained model after supervised fine-tuning) plays a fundamental role in RLHF to avoid overfitting. There is a line of theoretical RLHF work that modeled the problem as a reverse-KL regularized contextual bandit (Xiong et al., 2024a; Ye et al., 2024a; Zhong et al., 2024; Wu et al., 2024; Xie et al., 2024). However, they adopt the techniques from contextual bandits and neglect the power of reverse-KL-regularization, thus obtaining almost the same $\mathcal{O}(1/\epsilon^{2})$  [^1] For simplicity, we omit here the dependencies on quantities other than $\epsilon$. sample complexity as learning objectives without KL-regularization. Therefore, the question of
whether there exists a fundamental distinction between bandit learning objectives with and without KL-regularization
is still largely under-explored.
@@
Additionally, an emerging line of offline RLHF literature highlights the coverage of the reference policy $\pi_{0}$. The coverage of $\pi_{0}$ refers to the ability of the model to generate diverse responses for a wide range of prompts. A model with good coverage can generalize well to unseen contexts and actions, which is essential for the learned reward function to also generalize well. In practice, this is evidenced by the fact that the simple best-of-$n$ sampling based on $\pi_{0}$ is competitive with the well-tuned PPO algorithm for general open-ended conversation tasks (Dong et al., 2023), and the fact that the $\pi_{0}$ can solve a majority of the math problems with multiple responses (Shao et al., 2024; Nakano et al., 2021). While the coverage of $\pi_{0}$ is recognized as a key factor in offline RLHF, its impact on the sample complexity of online RLHF is still largely unknown. Thus, it is natural to ask:
If online RLHF is theoretically more efficient than offline RLHF under strong coverage of $\pi_{0}$?
In this paper, we answer the above questions by (1) providing a novel fine-grained decomposition for the suboptimiality of objective functions, which adapts to the strongly convex optimization landscape of the reverse-KL regularization and obtains a sharper sample complexity than the existing results, and (2) proposing an efficient 2-stage mixed sampling strategy for online RLHF with good coverage of $\pi_{0}$, which achieves sample complexity with only an additive dependence on the coverage coefficient. In contrast, the existing RLHF algorithms typically require a multiplicative dependence on the coverage coefficient.
@@ 1.1 Our Contributions
In this work, we make a first attempt to illustrate the statistical benefits of KL-regularization for contextual bandits and RLHF. Our main contributions are summarized as follows:
  • In Section 3, we study the contextual bandit problem with KL-regularization, which also serves as a mathematical formulation for RLHF with absolute-rating feedback. We provide a lower bound for the KL-regularized contextual bandit problem, which indicates that the sample complexity of the problem is $\Omega(\eta\log N_{\mathcal{R}}(\epsilon)/\epsilon)$ when $\epsilon$ is sufficiently small, where $N_{\mathcal{R}}(\epsilon)$ is the covering number of the reward function class and $\eta$ is the KL-regularization coefficient.
  • We povide a novel analysis to upper bound the suboptimality gap of the KL-regularized objective in contextual bandits, and propose a simple two-stage mixed sampling strategy to achieve a sample complexity of $\mathcal{O}(\max(\eta^{2}D^{2},\eta/\epsilon)\log N_{\mathcal{R}}(\epsilon/\delta))$ when the reward scale is a constant, where $D$ is the coverage coefficient of the reference policy $\pi_{0}$ and $\delta$ is the confidence parameter. To the best of our knowledge, this is the first work to provide an $\mathcal{O}(1/\epsilon)$ sample complexity for KL-regularized contextual bandits.
  • In Section 4, we extend our analysis to RLHF. We rigorously demonstrate that KL-regularization is essential for more efficient policy learning in RLHF with preference data. We further propose a two-stage mixed sampling strategy for online RLHF with good coverage of $\pi_{0}$, which achieves a sample complexity of $\mathcal{O}(\max(\eta^{2}D^{2},\eta/\epsilon)\log N_{\mathcal{R}}(\epsilon/\delta))$ when the reward scale is a constant.
@@ 1.2 Previous Understanding of KL Regularization in RL
Our analysis of KL-regularized contextual bandits and RLHF also contributes to the theoretical understanding of the impact of KL-regularization in RL since contextual bandits can be viewed as a simplified version of Markov decision processes (MDPs). In RL, KL-regularization has been widely used to stabilize the learning process and prevent the policy from deviating too far from the reference policy. Here, we provide a brief overview of the existing understanding of KL-regularization in decision-making problems. From the perspective of policy optimization, KL-regularization captures entropy regularization as a special case  [^2] We can regard the entropy regularization as a special case of KL-regularization by setting the reference policy as the uniform distribution., which is also an extensively used technique in RL literature (Sutton, 2018; Szepesvári, 2022). There is a large body of literature that has explored the benefits of entropy regularization or KL-regularization in RL (Schulman et al., 2015; Fox et al., 2016; Schulman et al., 2017a; Haarnoja et al., 2017, 2018; Ahmed et al., 2019). Most related to our work, Ahmed et al. (2019) provided a comprehensive understanding of the role of entropy regularization in RL, showing that entropy regularization can improve the training efficiency and stability of the policy optimization process by changing the optimization landscape through experiments on continuous control tasks (Brockman, 2016). Theoretically, Neu et al. (2017) provided a unified view of entropy regularization as approximate variants of Mirror Descent or Dual Averaging, and left the statistical justification for using entropy regularization in RL as an open question. Geist et al. (2019) provided a framework for analyzing the error propagation in regularized MDPs, which also focused on the proof of the convergence for the policy optimization methods with regularization and lacked a sharp sample complexity analysis.
2 Additional Related Work
Analyses of Policy Optimization with Regularization
While it is previously unknown whether regularization can improve the sample complexity of policy optimization without additional assumptions, there are some works that provided a sharp convergence rate in the presence of regularization (Mei et al., 2020; Shani et al., 2020; Agarwal et al., 2020, 2021; Lan, 2023). However, these works either assumed the access of exact or unbiased policy gradient or required uniform value function approximation error, which are not the standard case in general sample-based RL setting. For instance, Lan (2023) provided a sharp convergence rate for policy optimization with KL-regularization, assuming the access to an unbiased value function estimator (Condition 4.1) and the bounded infinity norm on the error (Conditions 4.2, 4.3), which is standard in the literature of optimization. However, RL algorithms usually make biased estimation to balance exploration and exploitation. Instead of focusing on the influence of regularization on the optimization, our work aims to understand how the KL-regularization affects the exploration and exploitation trade-off in the bandit and RLHF settings through a novel analysis on the optimal sample complexity.
@@ 2 Additional Related Work
There are mainly three types of RLHF algorithms: offline, online and hyrbid. The most well-known offline algorithms are Slic (Zhao et al., 2023b), Direct Preference Optimization (DPO) (Rafailov et al., 2024), Identity-PO (IPO) (Azar et al., 2024) and (SPIN) (Chen et al., 2024). They aim to approximate the closed-form solution of the optimization problem on a fixed offline dataset. For the online algorithms, the most representative one is Proximal Policy Optimization (PPO) (Schulman et al., 2017b). PPO has been used in the Chat-GPT (OpenAI, 2023), Gemini (Team et al., 2023), and Claude (Bai et al., 2022). However, the deep RL method PPO is known to be sample inefficient and unstable, making its success hard to reproduce for the open-source community. In response to this, there have been many efforts to propose alternative algorithms to the PPO algorithm. The Reward ranked fine-tuning (RAFT) (also known as rejection sampling finetuning) (Dong et al., 2023; Touvron et al., 2023; Gulcehre et al., 2023; Gui et al., 2024) is a stable framework requiring minimal hyper-parameter tuning, which iteratively learns from the best-of-n policy (Nakano et al., 2021). This framework proves to be particularly effective in the reasoning task such as (Gou et al., 2024; Tong et al., 2024). However, the RAFT-like algorithms only use the positive signal by imitating the best-of-n sampling. To further improve the efficiency, there is an emerging body of literature that proposes online direct preference optimization by extending DPO or IPO to an online iterative framework (Xiong et al., 2024a; Guo et al., 2024; Wu et al., 2024; Calandriello et al., 2024; Xiong et al., 2024b). Finally, for the third type, the common point of hybrid and online algorithms is that they both require further interaction with the preference oracle and on-policy data collection. The difference is that hybrid algorithms start with a pre-collected dataset (Xiong et al., 2024a; Song et al., 2024; Touvron et al., 2023), while the online algorithms learn from scratch.
@@
The theoretical study of RLHF can date back to the dueling bandits (Yue et al., 2012) and follow-up work on MDPs (Wang et al., 2023a; Zhu et al., 2023). However, these works deviate from the practice because they do not realize the significance of KL-regularization and still choose the greedy policy that simply maximizes the reward. After this line of work, Xiong et al. (2024a); Ye et al. (2024a); Song et al. (2024) highlight the KL-regularization theoretically and incorporate the KL term into the learning objective. However, they circumvent the special advantages of KL-regularization and still follow the techniques in bandit analysis, thus obtaining loose bounds. In our paper, we establish a new lower bound and a sharper upper bound for the KL-regularized framework, thus validating the empirical advantage of KL-regularization. There are also some works extending KL-regularized RLHF from bandit problems to the Markov decision process (MDP) problems (Zhong et al., 2024; Xiong et al., 2024b). We expect that our techniques can also be extended to the MDP setting, which we leave for future work.
3 KL-Regularized Contextual Bandits
In this section, we formally define the KL-regularized contextual bandit problem and provide a lower bound for the sample complexity of the problem. We then propose a novel two-stage mixed sampling strategy for online RLHF with good coverage of the reference policy $\pi_{0}$.
3.1 Problem Setup
In the contextual bandit setting, in each round $t$, the agent observes a context $x_{t}\in\mathcal{X}$ generated from a distribution $d_{0}$ and chooses an action $a_{t}\in\mathcal{A}$. The agent receives a stochastic reward $r_{t}\in\mathbb{R}$ depending on the context $x_{t}$ and the action $a_{t}$. The goal is to maximize the expected cumulative reward over $T$ rounds.